OpenStudy (anonymous):

Consider a triangle with one side of length a=9.8m and the angle opposite to side a is 60.0º. If the perimeter is 25.2m then find the largest angle of the triangle. Any ideas please?

OpenStudy (anonymous):

Does it mention what type of triangle? Or, is it just some randomly shaped triangle?

OpenStudy (anonymous):

no it doesnt but there is a figure of an oblique triangle

OpenStudy (anonymous):

What class is this for? The way I will answer this depends on it because I don't want to give you something too difficult.

OpenStudy (anonymous):

Geometry and linear algebra.. we also studied vector algebra

OpenStudy (anonymous):

Would it be possible to solve using the law of sin and cosin.. i did it but the angles i got only add up to 169.5 and im pretty sure my calculations were right

OpenStudy (anonymous):

That's what I was thinking.

OpenStudy (anonymous):

so how would that go?

OpenStudy (anonymous):

using sine Rule, \[{\sin60\over9.8}={\sin B\over25.2-9.8-c}={\sin C\over c}=2R\]

OpenStudy (anonymous):

\[ \sin B=\frac{\sqrt{3}\times(16.4-c)}{2\times9.8}\\ \sin C=\frac{c\sqrt{3}}{2\times9.8}\\ B+C=180-60=120^\circ \]